OpenStudy (anonymous):

Which of the following expressions does not simplify to a^2b+8b? (-4a^2b+8b) - ( -5a^b) (3a^2b+2b) -(2a^2b-6b) (2a^2b-4b^2) - (a^2b-4b^2+8b) (6a^2b+3a^2+7b) -(5a^2b-3a^2- b)

OpenStudy (anonymous):

I hate these kinds of problems. The only way to find the answer is to multiply them all out and simplify.

OpenStudy (anonymous):

*subtract and simplify.

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

Well, let's just begin with the first: (-4a^2b+8b) - (-5a^b) = -4a^2b + 8b + 5a^2b = a^2b + 8b This is good. (3a^2b+2b) - (2a^2b-6b) = 3a^2b + 2b - 2a^2b + 6b = a^2b + 8b This is good. (2a^2b-4b^2) - (a^2b-4b^2+8b) = 2a^2b - 4b^2 - a^2b + 4b^2 + 8b = a^2b + 8b This is good. I'm guessing this last one won't work: (6a^2b+3a^2+7b) -(5a^2b-3a^2- b) = 6a^2b + 3a^2 + 7b - 5a^2b + 3a^2 + b = a^2b + 6a^2 + 8b Yup, last one doesn't simplify.

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

you put a 2 in the last one

OpenStudy (anonymous):

I am sure is the first one

OpenStudy (anonymous):

it does not simplify? cus thts wut its askin

OpenStudy (anonymous):

Do you know how to simplify these KCB?

OpenStudy (anonymous):

wait both, first and last does't simplyfy

OpenStudy (anonymous):

doens't*

OpenStudy (anonymous):

yes kinda

OpenStudy (anonymous):

Show me how you would simplify the first one.